OpenStudy (anonymous):

how to integrate sin(t^3)dt?

OpenStudy (anonymous):

probably not possible

OpenStudy (anonymous):

i mean to find a closed form

OpenStudy (anonymous):

how so?

OpenStudy (anonymous):

just because you write down some function does not mean you can find another function in closed form whose derivative is that thing

OpenStudy (anonymous):

its a multiple choice question and no solution is not an option though

OpenStudy (anonymous):

you can always write \[\int_a^x\sin(t^3)dt\] as an anti - derivative, but you probably cannot find some function as an equation whose derivative you have

OpenStudy (anonymous):

what does the question say exactly?

OpenStudy (anonymous):

\[\int\limits_{x ^{2}}^{0} \sin(t ^{3}) dt\]sin(t^3)dt

OpenStudy (anonymous):

x2 will be at the top though and 0 will be at the bottom

OpenStudy (anonymous):

hold the phone. does it ask for the derivative???

OpenStudy (anonymous):

no just integrate

OpenStudy (anonymous):

find the derivative of \[\int_0^{x^3}\sin(t^3)dt\]? maybe?

OpenStudy (anonymous):

yes actually you are right im sorry

OpenStudy (anonymous):

how'd i guess?

OpenStudy (anonymous):

how would i solve that?

OpenStudy (anonymous):

the derivative of \[\int_0^{x^2}\sin(t^3)dt\] requires the chain rule . the derivative of the integral is the integrand, but you have a composite function because the upper limit of integration is x^2 not x

OpenStudy (anonymous):

so replace t by x^2 and then multiply by the derivative of x^2 which is 2x

OpenStudy (anonymous):

you will get \[\sin((x^2)^3)\times 2x=2x\sin(x^6)\]

OpenStudy (anonymous):

problem is actually very easy if you know what it is asking. the derivative of \[\int_a^xf(t)dt\] is \[f(x)\]

OpenStudy (anonymous):

but in your case you have a composition, it is not \[\int_a^xf(t)dt\] but rather \[\int_a^{x^2}f(t)dt\] so by the chain rule you replace t by x^2 and then multiply by the derivative

OpenStudy (anonymous):

for this i give myself the carnak award

OpenStudy (anonymous):

Just make a wild guess: Direvative of integral cancel out, then replace the limit = - sinx^6